A car is moving at a speed of 25 m/sec around a level curve of radius 100 m. What is the minimum value of the coefficient of static friction between the tires and the road required to prevent the car from skidding?


1
Expert's answer
2022-03-29T12:27:49-0400

The car will move without skidding when

"F=\frac{mv^2}{R}\leq\mu_s mg"

Hence, the minimum value of the coefficient of static friction between the tires and the road

"\mu_s=\frac{v^2}{gR}=\frac{25^2}{9.8*100}=0.64"